OpenStudy (anonymous):

solve: 5 sin^2(x) = cos x 0< x < 2

myininaya (myininaya):

remember sin^2x+cos^2x=1 so sin^2x=1-cos^2x so we have 5(1-cos^2x)=cosx 5-5cos^2x=cosx -5cos^2x-cosx+5=0 5cos^2x+cosx-5=0 this is a quadratic equation let u=cosx u^2=cos^2x so we have 5u^2+u-5=0 solve this for u and then don't foget u=cosx and then you can solve for x

OpenStudy (anonymous):

thank you